A normal person consumes 2000 food calories a day. If a food calorie is equivalent to 4200 J of energy,

use reasonable assumptions to calculate the number of kilograms of water that can be boiled away to steam starting from 20 celsius.

Show all your steps clearly including explanations for each step.
Answer #1

Answer -

given,

calorie consumption of normal person = 2000 food calories/day

1 food Calorie = 4200 J

Specific Heat of water = 4200 J/kgC [Known Value]

Initial Temperature of water = 20 C

Mass of water in kg that can be boiled = ?

We know that,

Heat = m * s * T ------------A

where,

m = mass

s = specific heat

T = change temperature

We know that,

Water boils at 100C. So, change in temperature T = 100C - 20 C = 80C

Now,

1 food Calorie = 4200 J

So,

2000 food calorie = 2000 * 4200 J = 8400000 J

So, Put the values in A

8400000 J = m * 4200 J/kgC * 80C

m = 8400000 J / (4200 J/kgC * 80C)

m = 25 kg

So, 25 kg of water can be boiled away to steam.